OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y A. (3, -1/2) B. (0, -2) C. (0, -11/2) D. (7, -5/2) I KNOW FOR SURE IT IS EITHER C OR B I WILL MEDAL AND FAN!

OpenStudy (anonymous):

@livias.random

OpenStudy (anonymous):

i am sure it is c

OpenStudy (anonymous):

@livias.random

OpenStudy (anonymous):

i just need u to check so i can have some comfort lol :)

OpenStudy (anonymous):

is i am wrong can u plz tell me how? thx

OpenStudy (anonymous):

if i am wrong i mean***

OpenStudy (anonymous):

x – 2 = – 2y x = -2y + 2 4(-2y + 2) + 2y = 11 -8y + 8 + 2y = 11 -6y + 8 = 11 -6y =3 y = -1/2 x = -2y + 2 x = -2(-1/2) + 2 x = 1 + 2 x = 3 (3,-1/2)

OpenStudy (anonymous):

oh snap

OpenStudy (anonymous):

So I got a, but I could have made a stupid mistake somewhere, do you want to show me your working?

OpenStudy (anonymous):

yes one sec let me type it

OpenStudy (anonymous):

x – 2 = – 2y x = -2y + 2 4(-2y + 2) + 2y = 11 -8y + 8 + 2y = 11 -6y + 8 = 11 -6y =3 y = -11/2 x = -2y + 3 x = -2(-1/2) + 2 x = 0+1 x = 0 OMG I KNOW WHAT I DID WRONG U WERE RIGHT!

OpenStudy (anonymous):

x – 2 = – 2y 0, -11/2 0 - 2 = -2 (-11/2) -2 = 22/2 -2 = 11 No - x – 2 = – 2y 3, -1/2 3 - 2 = -2 (-1/2) 1 = 2/2 1 = 1 Yes

OpenStudy (anonymous):

AWESOME

OpenStudy (anonymous):

ON THE SECOND PART I SAID -11/2 instead of -1/2

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

thx for cheking me!

OpenStudy (anonymous):

No worries :)